the distance between earth and moon (r) is 4.34 × 10⁸m.What is the magnitude of the gravitational force each exerts on the other?Earth mass (mE)= 5.97 × 1024 kg and Moon's mass (mM)=7.35 ×1022 kg


1
Expert's answer
2021-03-10T17:16:34-0500
"F=\dfrac{Gm_Em_M}{r^2},""F=\dfrac{6.67\cdot10^{-11}\ \dfrac{Nm^2}{kg^2}\cdot5.97\cdot10^{24}\ kg\cdot7.35\cdot10^{22}\ kg}{(4.34\cdot10^8\ m)^2},""F=1.55\cdot10^{20}\ N."